# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ IV.2.1 $\overline{B}\to D\overline{T}$ decay amplitudes
# Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\}})$
$\bar{1}^{\prime*}$ $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s}$ $C_{\bar{T}}$ $\bar{2}^{\prime*}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}C_{\bar{T}}$
$\bar{2}^{\prime*}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}C_{\bar{T}}$ $\bar{3}^{\prime*}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s}$ $C_{\bar{T}}$
$\bar{7}^{\prime*}$ $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}ss}$ $\sqrt{2}C_{\bar{T}}$ $\bar{8}^{\prime*}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}ss}$ $\sqrt{2}C_{\bar{T}}$
$\overline{13^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\prime-1}_{\bar{c}}$ $-\frac{\sqrt{3}}{2}C_{\bar{T}}$ $\overline{14^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\prime 0}_{\bar{c}}$ $-\frac{\sqrt{3}}{2}C_{\bar{T}}$
$\overline{15^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\prime-}_{\bar{c}s}$ $\frac{1}{2}C_{\bar{T}}$ $\overline{15^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\prime-}_{\bar{c}s}$ $\frac{1}{2}C_{\bar{T}}$
$\overline{15^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\prime-}_{\bar{c}s}$ $-C_{\bar{T}}$
$\overline{16^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\prime\prime-}_{\bar{c}}$ $\frac{1}{2}(C_{\bar{T}}+4E_{\bar{T}\bar{B}})$ $\overline{17^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\prime\prime 0}_{\bar{c}}$ $\frac{1}{2}(C_{\bar{T}}+4E_{\bar{T}\bar{B}})$
$\overline{18^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}(4T_{\bar{T}}+C_{\bar{T}})$ $\overline{18^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}(4T_{\bar{T}}+C_{\bar{T}})$
$\overline{18^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\prime\prime-}_{\bar{c}s}$ $2T_{\bar{T}}+C_{\bar{T}}+2E_{\bar{T}\bar{B}}$
# Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\}})$
$\bar{1}^{\prime*}$ $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s}$ $C^{\prime}_{\bar{T}}$ $\bar{2}^{\prime*}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}C^{\prime}_{\bar{T}}$
$\bar{4}^{\prime*}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}\bar{s}}$ $\sqrt{2}C^{\prime}_{\bar{T}}$ $\bar{5}^{\prime*}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}$ $C^{\prime}_{\bar{T}}$
$\bar{9}^{\prime*}$ $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}}$ $\sqrt{2}C^{\prime}_{\bar{T}}$ $\overline{10^{\prime}}^{*}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}}$ $\sqrt{\frac{2}{3}}C^{\prime}_{\bar{T}}$
$\overline{10^{\prime}}^{*}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}}$ $-\sqrt{\frac{2}{3}}C^{\prime}_{\bar{T}}$ $\overline{11^{\prime}}^{*}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}$ $-\sqrt{\frac{2}{3}}C^{\prime}_{\bar{T}}$
$\overline{13^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{2\sqrt{3}}C^{\prime}_{\bar{T}}$ $\overline{13^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{3}}C^{\prime}_{\bar{T}}$
$\overline{13^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime-}_{\bar{c}}$ $-\frac{\sqrt{3}}{2}C^{\prime}_{\bar{T}}$ $\overline{14^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\,\prime 0}_{\bar{c}}$ $\frac{1}{2\sqrt{3}}C^{\prime}_{\bar{T}}$
$\overline{15^{\prime}}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime-}_{\bar{c}s}$ $\frac{1}{2}C^{\prime}_{\bar{T}}$
$\overline{16^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{2}(4T^{\prime}_{\bar{T}}+C^{\prime}_{\bar{T}})$ $\overline{16^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $2T^{\prime}_{\bar{T}}+C^{\prime}_{\bar{T}}+2E^{\prime}_{\bar{T}\bar{B}}$
$\overline{16^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{2}(4T^{\prime}_{\bar{T}}+C^{\prime}_{\bar{T}})$ $\overline{17^{\prime}}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{2}(C^{\prime}_{\bar{T}}+4E^{\prime}_{\bar{T}\bar{B}})$
$\overline{18^{\prime}}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}(C^{\prime}_{\bar{T}}+4E^{\prime}_{\bar{T}\bar{B}})$
@@
From Table 7 we see that decay modes involving flavor exotic $\overline{T}$ are governed by internal $W$ amplitude $C_{\bar{T}}$, and, consequently, the sizes of their decay rates are highly related. Furthermore, as shown in Table 7 all twelve modes with $T$ in ${\bf 6}$ are highly related, as they are either proportional to $C_{\bar{T}}$ or $C^{\prime}_{\bar{T}}$ and we have $C^{\prime}_{\bar{T}}=\frac{V^{*}_{cd}}{V^{*}_{cs}}C_{\bar{T}}$. In particular, the following modes involving flavor exotic states have the following relations on rates,
$\Gamma(B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s})$ $=$ $\Gamma(\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s})=2\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s})=2\Gamma(B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s})$ (41)
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s})=2\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s})$
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}).$
@@
Note that all four flavor exotic states in scenario I are involved in these $\overline{B}\to D\overline{T}$ decays, but we need decay modes in $\Delta S=-1$ and $\Delta S=0$ transitions.
@@
Similarly all twenty four modes with $T$ in $\overline{\bf 15}$ are highly related, as they are either proportional to $C_{\bar{T}}$ or $C^{\prime}_{\bar{T}}$. In particular, for the following fourteen modes involving flavor exotic states in scenario II, we have
$2\Gamma(B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s})$ $=$ $4\Gamma(B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s})=4\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s})=2\Gamma(\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s})$ (42)
@@
$=$ $2\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s})=4\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s})$
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}\bar{s}})=2\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}})$
$=$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}})=3\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}})$
$=$ $3\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}})=3\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}).$
Note that even with both $\Delta S=-1$ and $\Delta S=0$ transitions, we only have nine out of twelve flavor exotic states involved in these $\overline{B}\to D\overline{T}$ decays. On the other hand, the above relations impose strong restrictions on these rates. The continue unobservation of some of these modes, such as $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}ss}$ and $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}ss}$ decays, will certainly pose tension to scenario II.
@@
$\Gamma(\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}})=\Gamma(\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}})=\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}),$ (43)
$\Gamma(B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}s})$ $=$ $\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s})=\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}})$ (44)

### v3

@@ IV.2.1 $\overline{B}\to D\overline{T}$ decay amplitudes
# Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ # Mode $A(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\}})$
$\bar{1}^{\prime*}$ $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s}$ $(C_{\bar{T}}+P_{\bar{T},1})$ $\bar{2}^{\prime*}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+P_{\bar{T},1})$
$\bar{2}^{\prime*}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}}+P_{\bar{T},1})$ $\bar{3}^{\prime*}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s}$ $(C_{\bar{T}}+P_{\bar{T},1})$
$\bar{7}^{\prime*}$ $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}ss}$ $\sqrt{2}(C_{\bar{T}}+P_{\bar{T},1})$ $\bar{8}^{\prime*}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}ss}$ $\sqrt{2}(C_{\bar{T}}+P_{\bar{T},1})$
$\overline{13^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\prime-1}_{\bar{c}}$ $\frac{\sqrt{3}}{2}(C_{\bar{T}}+P_{\bar{T},1})$ $\overline{14^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\prime 0}_{\bar{c}}$ $\frac{\sqrt{3}}{2}(C_{\bar{T}}+P_{\bar{T},1})$
$\overline{15^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\prime-}_{\bar{c}s}$ $\frac{1}{2}(C_{\bar{T}}+P_{\bar{T},1})$ $\overline{15^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\prime-}_{\bar{c}s}$ $\frac{1}{2}(C_{\bar{T}}+P_{\bar{T},1})$
$\overline{15^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\prime-}_{\bar{c}s}$ $(C_{\bar{T}}+P_{\bar{T},1})$
$\overline{16^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\prime\prime-}_{\bar{c}}$ $\frac{1}{2}((C_{\bar{T}}+P_{\bar{T},1})+4(E_{\bar{T}\bar{B}}+PA_{\bar{B}}))$ $\overline{17^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\prime\prime 0}_{\bar{c}}$ $\frac{1}{2}((C_{\bar{T}}+P_{\bar{T},1})+4(E_{\bar{T}\bar{B}}+PA_{\bar{B}}))$
$\overline{18^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}(4T_{\bar{T}}+P_{\bar{T},2}+(C_{\bar{T}}+P_{\bar{T},1}))$ $\overline{18^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\prime\prime-}_{\bar{c}s}$ $2(T_{\bar{T}}+P_{\bar{T},2})+(C_{\bar{T}}+P_{\bar{T},1})$
$\overline{18^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}(4T_{\bar{T}}+P_{\bar{T},2}+(C_{\bar{T}}+P_{\bar{T},1}))$ $2E_{\bar{T}\bar{B}}+2PA_{\bar{B}}$
# Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}[q^{\prime}q^{\prime\prime}]})$ # Mode $A^{\prime}(\overline{B}_{q}\to D\overline{T}_{\bar{c}\bar{q}\{q^{\prime}q^{\prime\prime}\}})$
$\bar{1}^{\prime*}$ $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s}$ $(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\bar{2}^{\prime*}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\bar{4}^{\prime*}$ $\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}\bar{s}}$ $\sqrt{2}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\bar{5}^{\prime*}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}$ $(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\bar{9}^{\prime*}$ $B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}}$ $\sqrt{2}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\overline{10^{\prime}}^{*}$ $B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\overline{10^{\prime}}^{*}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\overline{11^{\prime}}^{*}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\overline{13^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{2\sqrt{3}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\overline{13^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{3}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\overline{13^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{\sqrt{3}}{2}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$ $\overline{14^{\prime}}$ $\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\,\prime 0}_{\bar{c}}$ $\frac{1}{2\sqrt{3}}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\overline{15^{\prime}}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime-}_{\bar{c}s}$ $\frac{1}{2}(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$
$\overline{16^{\prime}}$ $B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{2}(4(T^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},2})+(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1}))$ $\overline{16^{\prime}}$ $\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $2(T^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},2})+(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})+2(E^{\prime}_{\bar{T}\bar{B}}+PE^{\prime}_{\bar{B}})$
$\overline{16^{\prime}}$ $\overline{B}{}_{s}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{2}(4(T^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},2})+(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1}))$ $\overline{17^{\prime}}$ $\overline{B}{}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}}$ $\frac{1}{2}((C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})+4(E^{\prime}_{\bar{T}\bar{B}}+PE^{\prime}_{\bar{B}}))$
$\overline{18^{\prime}}$ $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}((C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})+4(E^{\prime}_{\bar{T}\bar{B}}+PE^{\prime}_{\bar{B}}))$
@@
From Table 7 we see that decay modes involving flavor exotic $\overline{T}$ in scenario I are governed by $(C_{\bar{T}}+P_{\bar{T},1})$, and, consequently, their decay rates are highly related. Furthermore, as shown in Table 7, all twelve modes with $T$ in $\bf 6$} are highly related, as they are either proportional to $(C_{\bar{T}}+P_{\bar{T},1})$ or $(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$. In particular, the following modes involving flavor exotic states have relations on rates,
$\Gamma_{1}$ $\equiv$ $\Gamma(B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s})=\Gamma(\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s})=2\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s})=2\Gamma(B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s})$
$\Gamma_{2}$ $\equiv$ $\Gamma(B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s})=2\Gamma(\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s})=\Gamma(\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}}),$
$\Gamma_{1}$ $\simeq$ $\left|\frac{V_{cb}V^{*}_{cs}}{V_{cb}V^{*}_{cd}}\right|^{2}\Gamma_{2},$ (57)
@@
where the approximation in the last line neglects the penguin contributions. Note that all four flavor exotic states in scenario I are involved in these $\overline{B}\to D\overline{T}$ decays, but we need to consider decay modes in $\Delta S=-1$ and $\Delta S=0$ transitions.
One can roughly estimate the error in the approximation in Eq. (57) by using
$\frac{|P_{\bar{T},1}|}{|C_{\bar{T}}|}\sim\left|\frac{V_{tb}V^{*}_{ts}(c_{4}+\frac{c_{3}}{3})}{V_{cb}V^{*}_{cs}(c_{2}+\frac{c_{1}}{3})}\right|\simeq 0.19,$ (58)
and
$\frac{|P^{\prime}_{\bar{T},1}|}{|C^{\prime}_{\bar{T}}|}\sim\left|\frac{V_{tb}V^{*}_{td}(c_{4}+\frac{c_{3}}{3})}{V_{cb}V^{*}_{cd}(c_{2}+\frac{c_{1}}{3})}\right|\simeq 0.17,$ (59)
where $c_{i}$ are the Wilson coefficients, see Beneke:2001ev . Therefore, the errors of neglecting penguin contributions are roughly estimated to be of the order $\lambda$. To be more precise, we define the following ratio:
$R$ $\equiv$ $\left|\frac{V_{cb}V^{*}_{cd}}{V_{cb}V^{*}_{cs}}\right|^{2}\frac{\Gamma_{1}}{\Gamma_{2}}=\left|\frac{V_{cb}V^{*}_{cd}}{V_{cb}V^{*}_{cs}}\right|^{2}\frac{|A|^{2}+|\bar{A}|^{2}}{|A^{\prime}|^{2}+|\bar{A}^{\prime}|^{2}}$ (60)
$=$ $\left|\frac{V_{cb}V^{*}_{cd}}{V_{cb}V^{*}_{cs}}\right|^{2}\frac{|C_{\bar{T}}+P_{\bar{T},1}|^{2}+|\bar{C}_{\bar{T}}+\bar{P}_{\bar{T},1}|^{2}}{|C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1}|^{2}+|\bar{C}^{\prime}_{\bar{T}}+\bar{P}^{\prime}_{\bar{T},1}|^{2}},$
[…]
$R$ $=$ $\left|\frac{V_{cb}V^{*}_{cd}}{V_{cb}V^{*}_{cs}}\right|^{2}$
$\times\frac{|V_{cb}V^{*}_{cd}(c_{2}+\frac{c_{1}}{3})+\rho e^{i\phi}V_{tb}V^{*}_{ts}(c_{4}+\frac{c_{3}}{3})|^{2}+|V^{*}_{cb}V_{cd}(c_{2}+\frac{c_{1}}{3})+\rho e^{i\phi}V^{*}_{tb}V_{ts}(c_{4}+\frac{c_{3}}{3})|^{2}}{|V_{cb}V^{*}_{cd}(c_{2}+\frac{c_{1}}{3})+\rho e^{i\phi}V_{tb}V^{*}_{td}(c_{4}+\frac{c_{3}}{3})|^{2}+|V^{*}_{cb}V_{cd}(c_{2}+\frac{c_{1}}{3})+\rho e^{i\phi}V^{*}_{tb}V_{td}(c_{4}+\frac{c_{3}}{3})|^{2}},$
where we introduce a correction factor $\rho$ and a relative strong phase $\phi$ between the penguin and the tree amplitudes. We take the value of $\rho$ from 0 to 2 to parameterize our ignorance of the actual size of the penguin-to-tree amplitude ratio. By taking $\rho=0\sim 2$ and $\phi=0\sim 2\pi$, we obtain
$R-1=-0.20\sim 0.09.$ (62)
Hence, the correction on the approximation in the last line of Eq. (57) is estimated to be about 10 to 20% in the presence of penguin contributions. To effectively control the penguin pollution, further work is needed; see, for example, a study in a similar situation Fleischer:2007zn , where information of direct CP violation is needed. With these data, the sizes of the penguin and tree amplitudes, as well as the relative strong phase, can be probed.
@@
We turn to scenario II. Similarly all twenty four modes with $T$ in $\overline{\bf 15}$ are highly related, as they are either proportional to $(C_{\bar{T}}+P_{\bar{T},1})$ or $(C^{\prime}_{\bar{T}}+P^{\prime}_{\bar{T},1})$. In particular, for the following fourteen modes involving flavor exotic states in scenario II, we have
$\Gamma_{3}$ $\equiv$ $2\Gamma(B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}s})=4\Gamma(B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}s})=4\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}s})=2\Gamma(\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}s})$
@@
$\Gamma_{4}$ $\equiv$ $2\Gamma(B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}s})=4\Gamma(\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}s})=\Gamma(\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}\bar{s}})=2\Gamma(\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}\bar{s}})$
$=$ $\Gamma(B^{-}\to D^{+}\overline{T}^{\,--}_{\bar{c}})=3\Gamma(B^{-}\to D^{0}\overline{T}^{\,-}_{\bar{c}})=3\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,-}_{\bar{c}})=3\Gamma(\overline{B}{}^{0}\to D^{0}\overline{T}^{0}_{\bar{c}}),$
$\Gamma_{3}$ $\simeq$ $\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma_{4},$ (63)
where the approximation in the last line neglects the penguin contributions. Similarly, in the presence of penguin contributions, a correction of 10% to 20% on the above approximation is called for, see Eq. (62). Note that even with both $\Delta S=-1$ and $\Delta S=0$ transitions, we only have nine out of twelve flavor exotic states involved in these $\overline{B}\to D\overline{T}$ decays. On the other hand, the above relations impose strong restrictions on these rates. The continue unobservation of some of these modes, such as $B^{-}\to D_{s}^{+}\overline{T}^{\,--}_{\bar{c}ss}$ and $\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,-}_{\bar{c}ss}$ decays, will certainly pose tension to scenario II.
@@
$\Gamma(\overline{B}{}_{s}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}})=\Gamma(\overline{B}{}_{s}^{0}\to D^{0}\overline{T}^{\,\prime\prime 0}_{\bar{c}})\simeq\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(\overline{B}{}^{0}\to D_{s}^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s}),$ (64)
$\Gamma(B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}s})$ $=$ $\Gamma(\overline{B}{}^{0}\to D^{+}\overline{T}^{\,\prime\prime-}_{\bar{c}s})\simeq\left|\frac{V_{cs}}{V_{cd}}\right|^{2}\Gamma(B^{-}\to D^{0}\overline{T}^{\,\prime\prime-}_{\bar{c}})$ (65)